OpenStudy (anonymous):

Solve. |3 – 2x| = 9 A. x = –3 B. x = 6 C. x = –3, x = 6 D. x = 3, x = 6

zepdrix (zepdrix):

The absolute bars are telling us that there will be two solutions.\[\Large\rm |3 – 2x| = 9\]\[\Large\rm \pm(3 – 2x) = 9\]Rewrite the plus/minus on the other side, makes things easier to work with.

zepdrix (zepdrix):

\[\Large\rm 3 – 2x = \pm9\]

zepdrix (zepdrix):

Do you understand how to find your two solutions from this point?

OpenStudy (anonymous):

no:(

zepdrix (zepdrix):

So I guess we could start by subtracting 3 from each side, yes?

zepdrix (zepdrix):

\[\Large\rm -2x=\pm9-3\]

zepdrix (zepdrix):

Then to finish isolating the x we would divide both sides by -2,\[\Large\rm x=-\frac{1}{2}(\pm9-3)\]

zepdrix (zepdrix):

To find the first solution, let the 9 be positive, and simplify it down to a number. Then for your other solution, let the 9 be negative.

OpenStudy (anonymous):

I still don't really understand it though I mean i kinda do and kinda dont at the same time :/

OpenStudy (anonymous):

wait so are you saying add 9 to both side? because i'm not sure

zepdrix (zepdrix):

Ok let's try a different approach. I think maybe the plus/minus symbol is confusing you.\[\Large\rm |3-2x|=9\]We can drop the absolute bars by writing this as:\[\Large\rm 3-2x=9\]\[\Large\rm 3-2x=-9\]These two equations represent the absolute function. (They would normally be restricted to certain intervals of x, but we don't need to worry about that here).

zepdrix (zepdrix):

Do you know how to solve for x in this equation?\[\Large\rm 3-2x=9\]

zepdrix (zepdrix):

Or still confusing?

OpenStudy (anonymous):

yea I can do that

zepdrix (zepdrix):

That will give you your `first` solution.

zepdrix (zepdrix):

Solving the other equation for x,\[\Large\rm 3-2x=-9\]will give you your `second` solution.

OpenStudy (anonymous):

wait you add two to both sides right?

zepdrix (zepdrix):

No no no. Let's write it like this a sec,\[\Large\rm -2x+3=9\]3 is being `added` to each side, yes? So we do the inverse, `subtract` 3 from each side. -2 is `multiplying` the x, yes? so we do the inverse, `divide` both side by -2.

OpenStudy (anonymous):

ohhhhhhhhhhhhh ok thanks I think that I could do it now well atleast I think LOL but thx :)

zepdrix (zepdrix):

hehe